OpenStudy (anonymous):

The sum of two consecutive integers is at most 223. What is the larger of the two integers?

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (mrnood):

it can be any 2 integers (up to a certain level) because it tells you that the sum is AT MOST 223 for instance 1 & 2 2 & 3 18&19 all satisfy that inequality. In general if you have an interger x then the next integer is x+1 so the sum of 2 consecutive integers = 2x+1 you can use that to find the LARGEST possible number (but that is not what the question asks..)

OpenStudy (anonymous):

SO 3?

OpenStudy (mrnood):

no as i said for instanc 18&19 but also 100 & 101 READ what I said

OpenStudy (anonymous):

OK so what you just posted is 100 and 101.. The larger is 101

OpenStudy (mrnood):

you have missed the point What about 101 & 102 ? Have a little think before you post next

OpenStudy (texaschic101):

x + x+ 1 <= 223

OpenStudy (mrnood):

Did you type the question exactly as written?

OpenStudy (anonymous):

How am I missing the point?? The LARGER is the higher number....Therefore out of 102 its 103

OpenStudy (anonymous):

And yes I did!

OpenStudy (mrnood):

what about 103 & 104 have a think and read teh question

OpenStudy (texaschic101):

x + x + 1 <= 223 2x <= 222 x <= 111 x + 1 = 111 + 1 = 112

OpenStudy (texaschic101):

but the way the question is worded can be a bit confusing

OpenStudy (anonymous):

Very confusing!

OpenStudy (mrnood):

to be honest - this is not a well phrased question - it doesn't really have an answer as written. What @texaschic101 has done is given you the answer for a differnt question "what is the LARGEST possible value for one of the integers" but that is not what is asked

OpenStudy (texaschic101):

true ^^.....I think the way it is written, there can be more then 1 answer

OpenStudy (anonymous):

So this question is a tricky question or it didnt make it right

OpenStudy (texaschic101):

its tricky

OpenStudy (anonymous):

ahhhhhhhh!

OpenStudy (anonymous):

So what is the 2 answers you guys see?

OpenStudy (mrnood):

Please have a think and try to understand what we are saying it is not 2 answers - it is a NUMBER of answers ANY 2 integers less than 112 will answer teh question correctly. You need to look at the equation that was written above and understand HOW it relates to rthe question. doi not focus on the answer - this question does not have a proper answer

OpenStudy (anonymous):

I dont see an equation!

OpenStudy (mrnood):

IF the question was written properly then it would not be 'tricky' the some of integer x and the consecutive integer (x+1) is x+x+1 = 2x+1 THAT is the key point you needed to get to answer similar questions

OpenStudy (mrnood):

@texaschic101 wrote THIS eqution above x + x + 1 <= 223 2x <= 222 x <= 111 x + 1 = 111 + 1 = 112 BUT it doesn't answer your question, because your question is not posed properly

OpenStudy (anonymous):

Ok is there a way to solve this????

OpenStudy (mrnood):

I repeat: it is not 2 answers - it is a NUMBER of answers ANY 2 integers less than 112 will answer the question correctly.

OpenStudy (anonymous):

11,13,15 and 17

OpenStudy (anonymous):

what i found

OpenStudy (anonymous):

@MrNood

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

−112,−56,−28,−16,−14,−8,−7,−4,−2,−1,1,2,4,7,8,14,16,28,56,112

OpenStudy (anonymous):

Thats it

OpenStudy (texaschic101):

I am sorry buddy...I just don't know...there can be several answers

OpenStudy (anonymous):

:(